# Spatially covariant gravity with two degrees of freedom in the presence of an auxiliary scalar field: Hamiltonian analysis

### v1

@@ IV Degeneracy Analysis
To reduce the number of DOFs, additional conditions must be imposed on the Lagrangian (6). As we will see in this section, two conditions are required in order for the theory to propagate only two DOFs.
@@ IV.1 The first condition
As first-class constraints, $\pi_{i}$ and $\mathcal{C}_{i}$ correspond to spatial invariance. Without breaking spatial invariance, $\pi_{i}$ and $\mathcal{C}_{i}$ will always remain first-class constraints, which have vanishing Poisson brackets with all other constraints. For simplicity, we are allowed to omit these two first-class constraints in the subsequent degeneracy analysis. We therefore define a simplified Dirac matrix by omitting the two columns and rows corresponding to $\pi_{i}$ and $\mathcal{C}_{i}$:
@@
Our goal is to find conditions that eliminate one scalar-type degree of freedom. Generally, we have two options: one is to introduce extra conditions to convert second-class constraints into first-class constraints, and the other is to generate new secondary constraints. In both cases, a necessary condition is that the simplified Dirac matrix $\mathcal{M}^{ab}$ must be degenerate. The determinant of the simplified Dirac matrix $\mathcal{M}^{ab}$ is
@@
where the matrix $\mathcal{B}^{ab}$ is the lower-left corner of $\mathcal{M}^{ab}$. The degeneracy of $\mathcal{M}^{ab}$, i.e., $\det\mathcal{M}^{ab}=0$, thus implies that the sub-matrix $\mathcal{B}^{ab}$ of $\mathcal{M}^{ab}$ is degenerate, i.e., $\det\mathcal{B}^{ab}=0$.
At first glance, the determinant of the matrix $\mathcal{B}^{ab}$ may appear to be the matrix form of the first 2DOF condition. However, this is not the case. The 2DOF condition should be an equation satisfied by the action and should not contain conjugate momenta. Otherwise, we obtain a relation between the action and the conjugate momenta, which is in fact a constraint. Upon careful inspection, all elements of the matrix $\mathcal{B}^{ab}$ are independent of the conjugate momenta, except for $\mathcal{B}^{21}$, given by
@@
on the constraint surface of $\tilde{\pi}^{ij}$. In subsequent calculations, we will always use (37) as the value of $\mathcal{B}^{21}$ instead of (36). As a result, the matrix form of the first 2DOF condition becomes
@@
In the above, we have adjusted the coefficients of $\tilde{\pi}^{ij}$ and $\mathcal{C}$ in the matrix $\mathcal{B}^{ab}$ to make it more symmetric. Since $\mathcal{B}^{ab}$ is a degenerate matrix, there must exist a null eigenvector $\mathcal{V}_{a}$ satisfying
@@
In (41), $V(\vec{x})$ is an arbitrary function of the spatial coordinates, and $U_{a}(\vec{x},\vec{y})$ depends on the phase space variables. One method to obtain the explicit form of the null eigenvector $\mathcal{V}_{a}$ is to calculate the adjoint matrix of $\mathcal{B}^{ab}$, as detailed in Appendix C.
Assuming that $\mathcal{B}_{2}^{*}$ does not vanish, we can use a linear combination of the null eigenvector $\mathcal{V}_{a}$ and the constraints $\tilde{\pi}^{ij}$, $\mathcal{C}$, and $\frac{\delta S}{\delta\phi}$ to define a new constraint $\mathcal{C}^{\prime}$,
@@
In (42), $V(\vec{x})$ is the arbitrary function from (41).
The advantage of introducing the new constraint $\mathcal{C}^{\prime}$ is that its Poisson brackets with the constraints $p$, $\pi$, and $p^{kl}$ all vanish:
@@
With this new set of primary constraints, the consistency condition for the new constraint $\mathcal{C}^{\prime}$ becomes
$\frac{\mathrm{d}\mathcal{C}^{\prime}(\vec{x})}{\mathrm{d}t}$ $=\left[\mathcal{C}^{\prime}(\vec{x}),H_{\mathrm{T}}\right]$ (44)
@@
where $\lambda_{ij}$ has been fixed to zero in (26). If $\left[\mathcal{C}^{\prime}(\vec{x}),H_{\mathrm{C}}\right]\neq 0$, (44) yields a new secondary constraint $\Phi(\vec{y})$:
@@
Otherwise, if no secondary constraint arises, we can still find a new first-class constraint by taking a linear combination of the existing constraints $p$, $\pi$, and $p^{ij}$. The necessary and sufficient condition for generating secondary constraints is that the consistency condition is not automatically satisfied (instead of requiring the simplified Dirac matrix to be degenerate). Therefore, after imposing the consistency condition, two scenarios may arise:
@@
    The consistency condition is not automatically satisfied, and an additional secondary constraint is needed to enforce it;
@@
    The consistency condition is automatically satisfied, and a first-class constraint can be constructed from a linear combination of existing constraints.
When we impose a specific 2DOF condition, we must check whether it violates the consistency condition and leads to additional secondary constraints. Such a newly generated constraint must be taken into account in the subsequent analysis.
@@ IV.2 The second condition
where again $V(\vec{x})$ is the arbitrary function in (41). Similar to the constraint $\mathcal{C}^{\prime}$, the new constraint $\mathcal{\pi}^{\prime}$ has vanishing Poisson brackets:
@@
=$\mathcal{M}^{\prime}{}^{ab}=\begin{array}[]{|c|ccccc|}\hline\cr[\cdot,\cdot]&\mathcal{\pi}^{\prime}(\vec{y})&\left\{p(\vec{y}),p^{kl}(\vec{y})\right\}&\mathcal{C}^{\prime}(\vec{y})&\left\{\tilde{\pi}^{kl}(\vec{y}),\frac{\delta S}{\delta\phi(\vec{y})}\right\}&\Phi(\vec{x})\\ \hline\cr\mathcal{\pi}^{\prime}(\vec{x})&0&0&0&0&X\\ \left\{p(\vec{x}),p^{kl}(\vec{x})\right\}&0&0&0&-\mathcal{C}^{ba}(\vec{y},\vec{x})&X\\ \mathcal{C}^{\prime}(\vec{x})&0&0&X&X&X\\ \left\{\tilde{\pi}^{kl}(\vec{x}),\frac{\delta S}{\delta\phi(\vec{x})}\right\}&0&\mathcal{C}^{ab}(\vec{x},\vec{y})&X&X&X\\ \Phi(\vec{x})&X&X&X&X&X\\ \hline\cr\end{array}$ (48)
Since the first 2DOF condition $\mathcal{S}_{1}(\vec{x},\vec{y})$ has been imposed, which yields a secondary constraint that reduces the dimension of the phase space by one and thus removes half a degree of freedom, we need to impose an additional 2DOF condition to reduce the number of degrees of freedom to two. This means that the simplified Dirac matrix must be degenerate again. The determinant of this matrix is proportional to $\left[\mathcal{C}^{\prime}(\vec{x}),\mathcal{C}^{\prime}(\vec{y})\right]$, $\left[\Phi(\vec{x}),\mathcal{\pi}^{\prime}(\vec{y})\right]$, and $\det\mathcal{C}^{ab}(\vec{x},\vec{y})$. If the determinant of this matrix is required to vanish, two types of 2DOF conditions can be obtained:
@@
The 2DOF condition $\mathcal{S}_{2}^{(1)}$ will generate a secondary constraint if the consistency condition is not automatically satisfied. On the other hand, the 2DOF condition $\mathcal{S}_{2}^{(2)}$ clearly generates a first-class constraint.
If $\left[\mathcal{C}^{\prime}(\vec{x}),H_{\mathrm{C}}\right]=0$ in the previous section, the second 2DOF conditions become
@@
which are covered by the TTDOF conditions $\mathcal{S}_{2}^{(1)}$ and $\mathcal{S}_{2}^{(2)}$.
The space of theories satisfying the 2DOF conditions $\mathcal{S}_{1}$, $\mathcal{S}{}_{2}^{(1)}$, and $\mathcal{S}_{2}^{(2)}$ can be divided into three branches:
@@
    Linear dependence branch: If the secondary constraints are linearly dependent, the 2DOF conditions will be automatically satisfied. However, in this case, the number of degrees of freedom will not change. Due to the linear dependence, the number of secondary constraints is reduced by one, which offsets the half-degree-of-freedom reduction caused by the 2DOF conditions. For example, if the constraints $\tilde{\pi}^{ij}$, $\mathcal{C}$, and $\frac{\delta S}{\delta\phi}$ are linearly dependent, the 2DOF condition $\mathcal{S}_{1}$ is automatically satisfied. However, the number of degrees of freedom remains unchanged.
@@
    Non-physical branch: Some 2DOF conditions may render the theory trivial. For example, if the secondary constraints generated by the consistency condition still fail to satisfy the consistency condition, additional secondary constraints will be generated. These additional constraints may reduce the number of degrees of freedom to less than two, which is unphysical.
@@
    Physical branch: In this branch, we obtain a scalar-tensor theory with only two tensorial degrees of freedom.
In the following, when we refer to 2DOF conditions, we always mean those corresponding to the third physical branch.
@@ V.1 The Lagrangian
In [57], [58] (see also [74]), monomials of SCG have been constructed and classified according to the total number of derivatives $d$ in each monomial. As an example of our formalism, in this section we apply the two conditions to the concrete model with $d=2$. The action is given by
@@
As discussed earlier, we introduce the variable $B_{ij}$ to make the explicit inversion of the velocity $\dot{h}_{ij}$ in terms of the momentum possible. The equivalent action $S_{2}$ can be written as
@@
For later convenience, we evaluate some of the elements of $\mathcal{B}^{ab}$ as
@@
To prevent the dynamical terms in the theory from degenerating, we assume
@@
Other elements of the matrix $\mathcal{B}^{ab}$ are evaluated as
@@
where $\left(\frac{\delta^{2}S}{\delta\phi\delta\phi}\right)_{0}$ represents the terms in $\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta\phi(\vec{y})}$ that are independent of derivatives of the delta function. We use this notation because their explicit expressions are very complicated, and only the coefficients of the derivatives of the delta function are relevant to the subsequent analysis.
@@ V.2 The first condition
We have already obtained the first 2DOF condition in (38). To get the specific form of the 2DOF conditions for this model, a straightforward approach is to calculate the determinant of the matrix $\mathcal{B}^{ab}$, which is, however, very complicated and not illuminating. Alternatively, we can choose an equivalent approach by assuming that a null eigenvector $\mathcal{V}_{a}(\vec{x})$ exists and find the expression of this null eigenvector. In this approach, we can determine which conditions are required to make the Dirac matrix $\mathcal{B}^{ab}$ degenerate.
We assume $U_{a}(\vec{x},\vec{y})$ in (41) is independent of the derivative of the delta function. As a result, different orders of derivatives of the delta function will be separated in the matrix $\mathcal{B}^{ab}$,
@@
in which terms with different orders (of derivatives of the delta function) in the matrix $\mathcal{B}^{ab}$ separately satisfy
@@
Note that by observing the form of the secondary constraint $\mathcal{C}$ and $\frac{\delta S}{\delta\phi}$, (72) can only be weakly satisfied when $i=0$.
In the matrices $\mathcal{B}_{(1)}^{ab}(\vec{x},\vec{y})$ and $\mathcal{B}_{(2)}^{ab}(\vec{x},\vec{y})$, only the $\times 2$ 2 submatrices in the lower left corner are nonvanishing, which are
@@
respectively. In the above, a subscript “$x$” denotes functional dependence on $\vec{x}$. In the subsequent discussion, we omit $\vec{x}$ for brevity when the dependence on $\vec{x}$ is clear from the context.
According to (74), for the null eigenvalue equation of $\mathcal{B}_{(2)}^{ab}$, the part of matrix $\mathcal{B}^{ab}$ proportional to $\mathrm{D}^{x^{i}}\mathrm{D}_{x^{i}}\delta^{3}(\vec{x}-\vec{y})$ yields
$d_{2}\mathcal{V}_{2}+2Nd_{1}\mathcal{V}_{3}$ $=$ 0 0,0 , (75)
$2c_{4}\mathcal{V}_{2}+Nd_{2}\mathcal{V}_{3}$ $=$ 0 0.0 . (76)
As for the null eigenvalue equation of $\mathcal{B}_{(1)}^{ab}$, the part of matrix $\mathcal{B}^{ab}$ only related to $\mathrm{D}_{x^{i}}\delta^{3}(\vec{x}-\vec{y})$ yields
$\left(\frac{2d_{1}}{N}+2\frac{\partial d_{1}}{\partial N}-\frac{2}{N}\frac{\partial d_{2}}{\partial\phi}\right)\mathcal{V}_{2}-2\frac{\partial d_{1}}{\partial\phi}\mathcal{V}_{3}$ $=$ 0 0,0 , (77)
$\frac{2}{N}\frac{\partial c_{4}}{\partial\phi}\mathcal{V}_{2}-\left(2N\frac{\partial d_{1}}{\partial N}+2d_{1}\right)\mathcal{V}_{3}$ $=$ 0 0,0 , (78)
$\frac{2}{N^{2}}\frac{\partial c_{4}}{\partial\phi}\mathcal{V}_{2}-\left(2\frac{\partial d_{1}}{\partial N}+\frac{2d_{1}}{N}\right)\mathcal{V}_{3}$ $=$ 0 0,0 , (79)
$\left(\frac{2c_{4}}{N^{2}}-\frac{2}{N}\frac{\partial c_{4}}{\partial N}\right)\mathcal{V}_{2}-\left(\frac{2}{N}\frac{\partial c_{4}}{\partial\phi}-2\frac{\partial d_{2}}{\partial N}\right)\mathcal{V}_{3}$ $=$ 0 0.0 . (80)
There are four equations since the coefficients of $\mathrm{D}^{i}\phi$ and $a^{i}$ need to be proportional to the null eigenvector. We assume that $d_{1}$, $d_{2}$, and $c_{4}$ are nonvanishing. We can simply choose
@@
and obtain two independent equations
$d_{2}^{2}-4d_{1}c_{4}$ $=$ 0 0,0 , (83)
$\left(\frac{d_{1}}{N}+2\frac{\partial d_{1}}{\partial N}-\frac{2}{N}\frac{\partial d_{2}}{\partial\phi}\right)2Nd_{1}+\frac{\partial d_{1}}{\partial\phi}d_{2}$ $=$ 0 0.0 . (84)
The other two equations are not independent and can be obtained by combining these two.
@@
When we discuss the degeneracy of the matrix $\mathcal{B}_{(0)}^{ab}$, we need to be careful to identify which parts are weakly equal to zero. To clarify this, we first discuss how secondary constraints participate in the weak equality.
Since in the Lagrangian 58 the velocity term $\dot{h}_{ij}$ (encoded in $K_{ij}$) has been linearized, each canonical momentum corresponds to a primary constraint. Then all the secondary constraints can be combined into forms that are independent of canonical momenta. For example, we make a linear combination of the constraints $\tilde{\pi}^{ij}$ and $\mathcal{C}$,
@@
which does not depend on the special solution (85) chosen before.
Since we have replaced $\mathcal{B}^{21}$ with a term independent of the canonical momentum, (38) can be equivalently written as
@@
$\left(\mathcal{V}_{1}\right)_{ij}=-2d_{1}N\mathcal{V}_{ij}^{(N)}+d_{2}\mathcal{V}_{ij}^{(\phi)},$ (90)
@@
Since we have previously chosen a set of special solutions for $d_{1}$, $d_{2}$, and $c_{4}$, equation (72) with $i=0$ can be written as
@@
Since there are no terms such as $a_{i}a^{i}$ in equation (93), the coefficient $\beta(\vec{x},\vec{y})$ is fixed to zero. The above formula can equivalently be written as
@@
In order for (94) to have a solution, the coefficients of $B_{ij}B^{ij}$, $B^{2}$, and $R$ must vanish. There will be 6 equations for $c_{1}$, $c_{2}$, and $c_{3}$. We assume that
@@
This assumption simplifies the expression of $D_{2}^{b}$.

### v2

@@ IV Degeneracy Analysis
To reduce the number of DOFs, additional conditions must be put on the Lagrangian (6). As we will see in this section, two conditions are needed in order to make the theory propagate only two DOFs.
@@ IV.1 The first condition
As being the first-class constraints, $\pi_{i}$ and $\mathcal{C}_{i}$ correspond to the spatial invariance. Without breaking the spatial invariance, $\pi_{i}$ and $\mathcal{C}_{i}$ will always be kept as the first-class constraints, which have vanishing Poisson brackets with any other constraints. For the sake of simplicity, we are allowed to omit these two first-class constraints in the subsequent degeneracy analysis. We thus define a simplified Dirac matrix by omitting the two columns and rows corresponding to $\pi_{i}$ and $\mathcal{C}_{i}$:
@@
Our purpose is to find conditions to eliminate one scalar type degree of freedom. Generally we have two choices. One is to introduce extra conditions to change the second-class constraints into the first-class constraints, the other is to generate new secondary constraints. In both cases, a necessary conditions is that the simplified Dirac matrix $\mathcal{M}^{ab}$ must be degenerate. The determinant of the simplified Dirac matrix $\mathcal{M}^{ab}$ is
@@
where the matrix $\mathcal{B}^{ab}$ is the lower left corner of $\mathcal{M}^{ab}$. The degeneracy of $\mathcal{M}^{ab}$, i.e., $\det\mathcal{M}^{ab}=0$, thus implies that the sub-matrix $\mathcal{B}^{ab}$ of $\mathcal{M}^{ab}$ is degenerate, i.e., $\det\mathcal{B}^{ab}=0$.
Apparently, the determinant of matrix $\mathcal{B}^{ab}$ is the matrix form of the first 2DOF condition. This is not the case. Because the 2DOF condition should be an equation that the action satisfies, and should not contain conjugate momentum. Otherwise, we get the relationship between the action and the conjugate momentum, which is actually a constraint. By careful observation, all elements of the matrix $\mathcal{B}^{ab}$ are independent of conjugate momentum, except of $\mathcal{B}^{21}$ given by
@@
on the constraint surface of $\tilde{\pi}^{ij}$. In subsequent calculations, we always use (37) as the value of $\mathcal{B}^{21}$ instead of (36). As a result, the matrix form of the first 2DOF condition is
@@
In the above, we have adjusted the coefficients of $\tilde{\pi}^{ij}$ and $\mathcal{C}$ in the matrix $\mathcal{B}^{ab}$ to make the it more symmetric. Since $\mathcal{B}^{ab}$ is a degenerate matrix, there must be a null eigenvector $\mathcal{V}_{a}$ satisfying
@@
In (41), $V(\vec{x})$ is an arbitrary function of spatial coordinates and $U_{a}(\vec{x},\vec{y})$ depends on the phase space variables. One method to get the explicit form of the null eigenvector $\mathcal{V}_{a}$ is to calculate the adjoint matrix of matrix $\mathcal{B}^{ab}$, which can be found in Appendix C.
Assuming that $\mathcal{B}_{2}^{*}$ does not vanish, we can use the linear combination of the null eigenvector $\mathcal{V}_{a}$ and constraint $\tilde{\pi}^{ij}$, $\mathcal{C}$, $\frac{\delta S}{\delta\phi}$ to define a new constraint $\mathcal{C}^{\prime}$,
@@
In (42), $V(\vec{x})$ is the arbitrary function in (41).
The merit of introducing the new constraint $\mathcal{C}^{\prime}$ is that the Poisson brackets between $\mathcal{C}^{\prime}$ and constraints $p$, $\pi$, $p^{kl}$ are all vanishing
@@
With this new set of primary constraints, the consistency conditions of the new constraint $\mathcal{C}^{\prime}$ reduce to
$\frac{\mathrm{d}\mathcal{C}^{\prime}(\vec{x})}{\mathrm{d}t}$ $=\left[\mathcal{C}^{\prime}(\vec{x})(\vec{x}),H_{\mathrm{T}}\right]$ (44)
@@
where $\lambda_{ij}$ has been fixed to be zero in (26). If $\left[\mathcal{C}^{\prime}(\vec{x}),H_{\mathrm{C}}\right]\neq 0$, (44) yields a new secondary constraint $\Phi(\vec{y})$:
@@
Otherwise, there is no secondary constraint, but we can always find a new first class constraint through linear combination of constraints $p$, $\pi$ and $p^{ij}$. Because the necessary and sufficient condition for generating secondary constraints is that the consistency condition cannot be automatically satisfied, rather than requiring the simplified Dirac matrix to be degenerate. Therefore, after requiring the consistency condition, two cases may arise:
@@
    The consistency condition is not automatically satisfied, so that an additional secondary constraint is required to ensure that the consistency condition is satisfied;
@@
    The consistency condition is automatically satisfied, and a first class constraint is obtained by combining the existing constraints.
When we require a specific 2DOF condition, we should check whether this condition violates the consistency condition and generates additional secondary constraints. We should consider this newly generated secondary constraint in the subsequent discussion.
@@ IV.2 The second condition
where again $V(\vec{x})$ is the arbitrary function in (41). Similar to constraint $\mathcal{C}^{\prime}$, the new constraint $\mathcal{\pi}^{\prime}$ has the vanishing Poisson brackets
@@
[…]
Since the first 2DOF condition $\mathcal{S}_{1}(\vec{x},\vec{y})$ has been imposed, which yields a secondary constraint to reduced the dimension of the phase space by one and thus reduce a half degree of freedom. We need to require an additional 2DOF condition to reduce the number of degrees of freedom to 2, which means that the simplified Dirac matrix is required to be degenerate again. The determinant of the matrix is proportional to $\left[\mathcal{C}^{\prime}(\vec{x}),\mathcal{C}^{\prime}(\vec{y})\right]$, $\left[\Phi(\vec{x}),\mathcal{\pi}^{\prime}(\vec{y})\right]$ and $\det\mathcal{C}^{ab}(\vec{x},\vec{y})$. If the determinant of this matrix is required to be vanishing, two types of 2DOF conditions can be obtained:
@@
The 2DOF condition $\mathcal{S}_{2}^{(1)}$ will generate a secondary constraint, if the consistency condition is not automatically satisfied. On the other hand, the 2DOF condition $\mathcal{S}_{2}^{(2)}$ clearly generates a first-class constraint.
If $\left[\mathcal{C}^{\prime}(\vec{x}),H_{\mathrm{C}}\right]=0$ in the previous section, the second 2DOF conditions will become
@@
which is covered by TTDOF conditions $\mathcal{S}_{2}^{(1)}$and $\mathcal{S}_{2}^{(2)}$.
The space of theories satisfying the 2DOF conditions $\mathcal{S}_{1}$, $\mathcal{S}{}_{2}^{(1)}$ and $\mathcal{S}_{2}^{(2)}$ can be divided into three branches:
@@
    Linear constraints branch: If the secondary constraints are linearly dependent, 2DOF conditions will be strongly vanishing. However, in this case the number of degrees of freedom will not change. Due to the linear dependence, the number of secondary constraints will be reduced by one, which will offset the half degree of freedom reduction caused by the 2DOF conditions. For example, if constraints $\tilde{\pi}^{ij}$, $\mathcal{C}$ and $\frac{\delta S}{\delta\phi}$ are linearly dependent, 2DOF condition $\mathcal{S}_{1}$ is automatically satisfied. However, in this case the number of degrees of freedom remains unchanged.
@@
    Non-physical branch: Some 2DOF conditions may make the theory be trival. For example, if the secondary constraints generated by the consistency condition still cannot satisfy the consistency condition, additional secondary constraints will be generated. However, these additional secondary constraints may reduce the number of degrees of freedom to be less than 2, which is an unphysical situation.
@@
    Physical branch: In this branch, we get a scalar-tensor theory with only two tensorial degrees of freedom.
In the following when we mention 2DOF conditions, we always refer to the third physical branch.
@@ V.1 The Lagrangian
In [57, 58] (see also [74]), monomials of SCG have been constructed and classified according to the total number of derivatives $d$ in each monomial. As an example of our formalism, in this section we will apply the two conditions to the concrete model of $d=2$. The action of is given by
@@
As being discussed before, we introduce the variable $B_{ij}$ in order to make the explicit reversion of velocity $\dot{h}_{ij}$ in terms of the momentum possible. The equivalent action $S_{2}$ can be written as
@@
For later convenience, we can evaluate some of the elements of $\mathcal{B}^{ab}$ as
@@
In order to prevent the dynamical terms of the theory from degenerating, we assume
@@
Other elements of matrix $\mathcal{B}^{ab}$ are evaluated to be
@@
where $\left(\frac{\delta^{2}S}{\delta\phi\delta\phi}\right)_{0}$ represents terms in $\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta\phi(\vec{y})}$ that are independent of derivatives of the delta function. We use this notation because their specific expressions are very complicated and only the coefficients of derivatives of the delta function are relevant to the subsequent analysis.
@@ V.2 The first condition
We have already obtained the first 2DOF condition in (38). To get the specific form of the 2DOF conditions for the this model, a straightforward approach is to calculate the determinant of the matrix $\mathcal{B}^{ab}$, which is however is very complicated and not illuminating. Alternatively, we can choose an equivalent approach by assuming that a null eigenvector $\mathcal{V}_{a}(\vec{x})$ exists and find the expression of this null eigenvector. In this approach, we can find which conditions are required to make the Dirac matrix $\mathcal{B}^{ab}$ degenerate.
We assume $U_{a}(\vec{x},\vec{y})$ in (41) is independent of the derivative of the delta function. As a result, different orders of derivative of the delta function will be separated in the matrix $\mathcal{B}^{ab}$,
@@
in which terms with different order (of derivative of the delta function) in the matrix $\mathcal{B}^{ab}$ separately satisfies
@@
Note by observing the form of the secondary constraint $\mathcal{C}$ and $\frac{\delta S}{\delta\phi}$, (72) can only be weakly satisfied when $i=0$.
In matrix $\mathcal{B}_{(1)}^{ab}(\vec{x},\vec{y})$ and $\mathcal{B}_{(2)}^{ab}(\vec{x},\vec{y})$, only the $2\times 2$ sub-matrices in the lower left corner are not vanishing, which are
@@
respectively. In the above a subscript “xx” denotes functional dependence on $\vec{x}$. In the subsequent discussion, we omit $\vec{x}$ for brevity when the dependence on $\vec{x}$ is clear from the context.
According to (74), for the null eigenvalue equation of $\mathcal{B}_{(2)}^{ab}$, the part of matrix $\mathcal{B}^{ab}$ is proportional to $\mathrm{D}^{x^{i}}\mathrm{D}_{x^{i}}\delta^{3}(\vec{x}-\vec{y})$, which yields
$d_{2}\mathcal{V}_{2}+2Nd_{1}\mathcal{V}_{3}$ $=$ $0,$ (75)
$2c_{4}\mathcal{V}_{2}+Nd_{2}\mathcal{V}_{3}$ $=$ $0.$ (76)
As for the null eigenvalue equation of $\mathcal{B}_{(1)}^{ab}$, the part of matrix $\mathcal{B}^{ab}$ only related to $\mathrm{D}_{x^{i}}\delta^{3}(\vec{x}-\vec{y})$, which yields
$\left(\frac{2d_{1}}{N}+2\frac{\partial d_{1}}{\partial N}-\frac{2}{N}\frac{\partial d_{2}}{\partial\phi}\right)\mathcal{V}_{2}-2\frac{\partial d_{1}}{\partial\phi}\mathcal{V}_{3}$ $=$ $0,$ (77)
$\frac{2}{N}\frac{\partial c_{4}}{\partial\phi}\mathcal{V}_{2}-\left(2N\frac{\partial d_{1}}{\partial N}+2d_{1}\right)\mathcal{V}_{3}$ $=$ $0,$ (78)
$\frac{2}{N^{2}}\frac{\partial c_{4}}{\partial\phi}\mathcal{V}_{2}-\left(2\frac{\partial d_{1}}{\partial N}+\frac{2d_{1}}{N}\right)\mathcal{V}_{3}$ $=$ $0,$ (79)
$\left(\frac{2c_{4}}{N^{2}}-\frac{2}{N}\frac{\partial c_{4}}{\partial N}\right)\mathcal{V}_{2}-\left(\frac{2}{N}\frac{\partial c_{4}}{\partial\phi}-2\frac{\partial d_{2}}{\partial N}\right)\mathcal{V}_{3}$ $=$ $0.$ (80)
There are four equations since the coefficients of $\mathrm{D}^{i}\phi$ and $a^{i}$ need to be proportional to the null eigenvector. We assume that $d_{1}$, $d_{2}$ and $c_{4}$ are not vanishing. We can simply choose
@@
and we can get two independent equations
$d_{2}^{2}-4d_{1}c_{4}$ $=$ $0,$ (83)
$\left(\frac{d_{1}}{N}+2\frac{\partial d_{1}}{\partial N}-\frac{2}{N}\frac{\partial d_{2}}{\partial\phi}\right)2Nd_{1}+\frac{\partial d_{1}}{\partial\phi}d_{2}$ $=$ $0.$ (84)
The other two equations are not independent and can be obtained by combining these two equations.
@@
When we discuss the degeneracy of the matrix $\mathcal{B}_{(0)}^{ab}$, we need to be careful to identify which parts weakly equal to zero. To make this clear, we should first discuss how secondary constraints participate in the weakly equal equation.
Since in the Lagrangian 58 the the velocity term $\dot{h}_{ij}$ (encoded in $K_{ij}$) has been linearized, each canonical momentum will correspond to a primary constraint. Then all the secondary constraints can be combined into a form that are independent of canonical momenta. For example, we make a linear combination of the constraints $\tilde{\pi}^{ij}$ and $\mathcal{C}$,
@@
which does not depend on the special solution (85) we chose before.
Since we have replaced $\mathcal{B}^{21}$ with a term which is independent of the canonical momentum, (38) can be equivalently written as
@@
$\left(\mathcal{\mathcal{V}}_{1}\right)_{ij}=-2d_{1}N\mathcal{V}_{ij}^{(N)}+d_{2}\mathcal{V}_{ij}^{(\phi)},$ (90)
@@
Since we have previously chosen a set of special solutions for $d_{1}$, $d_{2}$ and $c_{4}$, equation (72) with $i=0$ can be written as
@@
Since there are no terms such as $a_{i}a^{i}$ in equation (93), the coefficient $\beta(\vec{x},\vec{y})$ is fixed to zero. The above formula can be equivalently written as
@@
In order to make (94) have a solution, the coefficients of $B_{ij}B^{ij}$, $B^{2}$ and $R$ need to be vanishing. There will be 6 equations for $c_{1}$, $c_{2}$ and $c_{3}$. We assume that
@@
This assumption will simplify the expression of $D_{2}^{b}$.